# Spend Wisely: Maximizing Post-Training Gains in Iterative Synthetic Data Boostrapping

### v1

@@ 4 Experiments
We conduct three experiments to evaluate our schemes: the synthetic setting with Gaussian data as in Section 3.2, image denoising with diffusion probabilistic models, and mathematical reasoning with LLMs. These experiments cover diverse settings, including a toy scenario to validate the theory and two large-scale tasks in image and natural language processing. The setups and implementation details are summarized in Table 1.
In the two real-world settings, we emphasize that our goal is not to propose a novel task-solving method or generation technique. Rather, we focus on optimizing the allocation of generation and training costs in iterative learning with synthetic data—an objective that is method-agnostic and can be applied to any generation approach. Our aim here is to determine the optimal generation strategy $n_{t}$ given a predefined synthetic generation method.
4.1 Synthetic Experiment with Gaussian Distributions
@@
We begin with the synthetic setting introduced in our warm-up (Section 3.2), which involves optimizing a generative distribution parameterized as a Gaussian to maximize the expected reward. Specifically, we focus on a two-dimensional space using MLE updates at each iteration. The generator is set to $\mathcal{N}(\bm{\theta},\bm{I}_{2})$, with fixed variance. The reward model is defined as an exponential function $\mathcal{R}(\bm{x})=\exp\left(-\lVert\bm{x}\rVert^{2}/(2\kappa^{2})\right)$, favoring data concentrated near the origin. $\kappa$ controls the flatness of the reward.
All theoretical results from the warm-up extend naturally to the high-dimensional case, as detailed in Section E.1. We now validate that with equal computational budgets and iterations, the exponential scheme outperforms the linear scheme, which outperforms the constant scheme.
Implementation.
We consider two settings with $\kappa^{2}=2,4$. Based on theoretical analysis, the exponential scheme is set as $n_{t}=\lfloor n_{0}\cdot(1+1/\kappa^{2})^{t}\rfloor$, where $n_{0}=10$. For a fixed total budget and number of iterations, the constant scheme is then $n_{t}=\left\lfloor T^{-1}\sum_{k=0}^{T-1}n_{0}\cdot(1+1/\kappa^{2})^{k}\right\rfloor$, while the linear scheme is $n_{t}=\left\lfloor\frac{2(t+1)}{T(T-1)}\sum_{k=0}^{T-1}n_{0}(1+1/\kappa^{2})^{k}\right\rfloor$. The initial parameter is set to $\bm{\theta}^{(0)}=(1,1)^{\top}$, and the optimal solution for the reward function is $\bm{\theta}^{*}=(0,0)^{\top}$. The optimal reward value is then $\mathbb{E}_{\bm{x}\sim\mathcal{N}(\bm{0},\bm{I})}[\mathcal{R}(\bm{x})]=(1+1/\kappa^{2})^{-1}.$ We evaluate how fast the three schemes approach this optimal reward value.
Figure 2: Empirical results of the toy example. We compare the exponential, constant, and linear policy, and show the gap to the optimal expected reward as a function of the computational cost ($\sum_{t}n_{t}$, consistent with the setup in Section 3.2). Averaged over 1,000 runs.
@@ Results.
We have repeated the training 1,000 times for each scheme and average the performance. Figure 2 plots the gap to the optimal reward as a function of the cumulative cost, defined as the total computational cost incurred up to that point. Additional empirical results with diverse parameters are left in Section F.1. The results demonstrate that the exponential policy consistently approaches the optimal reward faster than the other schemes, reducing the cost to achieve the same performance by up to 60% compared to the constant policy. While the linear policy converges slightly more slowly, the constant policy, as shown in the left figure, clearly fails to converge. These findings align well with our theoretical proof that the exponential scheme is optimal (Theorem 3.1) and that the constant policy cannot converge to optimal with high probability (Theorem 3.2).
4.2 Image Denoising with Diffusion Models
Iterative learning with synthetic data is widely used in image processing, particularly for adapting trained models to new distributions where annotated or paired data are scarce, but unannotated data are abundant (Kirillov et al., 2023; Hemmat et al., 2024; Azizi et al., 2023; Fan et al., 2024). In these scenarios, the trained model generates synthetic annotations for the unannotated data. These generated pairs are filtered and subsequently used to fine-tune the model. This iterative process enables better adaptation to the target distribution and is often repeated across multiple iterations.
Figure 3: Empirical results of image denoising. The figure shows the average PSNR of the generated denoised images as a function of computational cost. Computational cost is measured in floating-point operations (FLOPs) during iterative learning, including both generation and training. The training batch size is set to $B=640$. We adopt $n_{t}=1.1^{t}\cdot B$ and $n_{t}=1.05^{t}\cdot B$ as the exponential growth policies for $s=0.01$ and $s=0.02$, respectively.
Setup.
In line with these works, we apply iterative learning to the image denoising task, a representative problem in image processing, where the goal is to recover $\bm{x}$ from a noisy observation $\bm{y}$,
$\bm{y}=\bm{x}+\bm{\eta},\quad\bm{\eta}\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I}).$
Here, $\bm{\eta}$ represents an unknown Gaussian noise with fixed variance $\sigma^{2}$. We choose diffusion probabilistic models (DPMs) (Ho et al., 2020; Song et al., 2021) as denoisers due to their intrinsic design, which aligns closely with the denoising process. Specifically, the forward process of a DPM adds Gaussian noise to a clean image, while the reverse process gradually estimates and removes this noise using a trained network, effectively performing denoising. Readers may refer to Appendix D for more details. In our setting, the observed $\bm{y}$ as a Gaussian perturbation of $\bm{x}$ follows the same distribution as an intermediate step of the diffusion process up to a scalar (Xie et al., 2023). For a known noise variance $\sigma^{2}$, we can match it with a certain intermediate step $s$ in the diffusion process. With the match, we apply the reverse denoising steps of the DPM to iteratively remove noise, taking the final output as the denoised image.
A key challenge in image denoising is that we only have access to noisy observations and lack their corresponding clean counterparts. We use the iterative learning framework to annotate synthetic ‘clean’ images to improve the denoising. In each iteration, the diffusion model first generates synthetic denoised images $\hat{\bm{x}}$ from the noisy observations. These are then filtered with a reward model using the Peak Signal-to-Noise Ratio (PSNR) values compared to the real images $\bm{x}$. The reward is defined as
$\mathcal{R}(\hat{\bm{x}},\bm{x})=\frac{\text{PSNR}(\hat{\bm{x}},\bm{x})-r_{\text{min}}}{r_{\text{max}}-r_{\text{min}}}.$
Notably, we only allow the reward model to have access to the ground truth $\bm{x}$. The filtered synthetic data are then used to fine-tune the model through a standard noise-prediction training procedure, iteratively improving its denoising capability.
Table 2: Best PSNR of each policy on image denoising. The final PSNR values are determined using a held-out validation set. We report all the policies in Figure 3. We use boldface for the best result while the underline for the second best.
Policy $n_{t}$ $s=0.01$ $s=0.02$
@@ 4.2 Image Denoising with Diffusion Models
exponential - 44.84 40.14
Implementations.
We use an unconditional generative model [^1] https://huggingface.co/google/ddpm-cifar10-32 trained on CIFAR10 (Krizhevsky, 2009) and fine-tune it to perform denoising on the MNIST (Deng, 2012) dataset. We perform denoising for two noise levels corresponding to the diffusion step $s\in\{0.01,0.02\}$. A summary of the algorithm is provided in Algorithm 2.
We consider three constant and three linear policies, each with small, medium, and large $n_{0}$. For the exponential policy, we have performed a small hyperparameter search over $n_{0}$ and $\mu$, selecting good configurations for two settings. The complete results of the hyperparameter search are presented as an ablation study on different exponential schemes in Section F.2, demonstrating its robustness. The configurations for all schemes are shown in Figure 3. We include the performance of the pre-trained model without fine-tuning for comparisons.
Figure 4: Empirical results of math reasoning. The figure shows the accuracies with respect to the computational cost, measured in FLOPs as well. The training batch size is set to $B=256$. We adopt $n_{t}=10\cdot 2^{t}\cdot B$ as the exponential policy. We set the temperature to 0.3 during the generation step and include the results for a temperature of 0.7 in Section F.4, which lead to the same conclusions.
Results.
We summarize the final performance of different schemes and baselines in Table 2 and show the trade-off between performance and computational cost in Figure 3. We observe that the exponential policy yields the best performance for $s=0.01$, and performs comparable to the best linear policy for $s=0.02$. This observation demonstrates both the benefits of an increasing policy (Theorem 3.3) and the robustness of the exponential growth policy (Theorem 3.5). In contrast, the constant policy only matches the exponential growth policy in the early stages of training, but soon reaches a performance plateau and eventually decays due to overfitting, corroborating Theorem 3.2.
4.3 Math Reasoning with LLMs
Now we move to natural language processing. Iterative learning with synthetic data is also an emerging method for LLM. Previous research (Zelikman et al., 2022) leverages the model to generate synthetic data, which is then filtered using a reward model or accuracy metrics and subsequently utilized to fine-tune the model. This framework is employed in various settings, including instruction tuning (Touvron et al., 2023), alignment (Dong et al., 2023), and recent efforts to enhance reasoning capabilities through chain-of-thought (CoT) approaches (Hosseini et al., 2024; Lehnert et al., 2024; Singh et al., 2025).
Setup.
We select mathematical reasoning to evaluate the effectiveness of our strategy in iterative learning. Let $q$ and $a$ represent the question and its corresponding answer, respectively. We use the LLM to generate synthetic solutions with CoT and in-context examples. These synthetic data are filtered based on the correctness of the final generated answer $\hat{a}$, defined as the reward $\mathcal{R}(\hat{a})=\mathbf{1}(\hat{a}=a).$ The filtered data are used to further train the model in an iterative process, without filtering based on the correctness of CoT itself. In summary, synthetic data with correct final solutions are generated and used for training in each iteration, closely mimicking the STaR method (Zelikman et al., 2022).
Implementations.
Many contemporary LLMs are known to overfit to GSM8k (Cobbe et al., 2021) or exhibit a certain degree of data leakage (Zhang et al., 2024; Mirzadeh et al., 2024), where even minor changes to numbers or names can result in significant performance degradation. This issue renders the evaluation results on GSM8k unreliable. To address this, we regenerate three math datasets—GSM_symbolic, GSM_p1, and GSM_p2—using symbolic math templates from Mirzadeh et al. (2024), a highly cited paper on evaluating math reasoning. Each dataset consists of questions paired with their correct answers. GSM_symbolic retains a similar difficulty level to GSM8k, while GSM_p1 and GSM_p2 introduce one and two additional clauses per question, respectively, making them progressively more challenging. Together, these datasets complete a robust evaluation set.
A summary of the algorithm is provided in Algorithm 3. Similarly to image denoising, we consider three constant and three linear policies, each with small, medium, and large $n_{0}$. For the exponential scheme, we have selected one configuration and left the full ablation results in Section F.3. The configurations for all schemes are detailed in Figure 4 and its caption. We also include the performance of the pre-trained model as a baseline.
Table 3: Best accuracy of each policy on math reasoning. The final accuracy are determined using a held-out validation set. We report all the policies in Figure 4.
Policy $n_{t}$ symbolic p1 p2
Pre-trained N/A 55.60 34.98 15.72
constant $10\cdot B$ 60.28 38.50 16.93
@@ 4.3 Math Reasoning with LLMs
exponential - 65.66 47.26 20.65
@@ Results.
Figure 4 shows the performance versus computational cost across the three datasets under iterative training, while Table 3 summarizes the final performance for all policies. Across all three difficulty levels, the exponential growth policy demonstrates the steadiest improvement compared to other schemes. Although the linear policy ranks second overall, it fails to match the exponential growth policy when the training budget is large. In particular, with the exponential growth policy, iterative learning increases accuracy on GSM_p1 from 35% to 47% and on GSM_p2 from 16% to 21%. Based on these results, we recommend the exponential scheme for iterative learning.
@@ 5 Conclusion
In this work, we take the first step toward understanding how to design optimal budget allocation policies in iterative bootstrapping for post-training. Through a combination of theoretical analysis and empirical validations, we demonstrate that constant policies are insufficient for achieving convergence, whereas increasing policies offer a more effective alternative. Among increasing policies, the exponential growth policy emerges as a robust and efficient method.
@@
Impact Statement
This research advances the field of machine learning by addressing the strategic allocation of computational resources in the post-training phase using synthetic data. We introduce a novel theoretical framework to optimize the distribution of a fixed budget across multiple training iterations, an approach that demonstrates significant improvement in model performance. Our findings challenge the conventional constant allocation strategies, revealing that dynamically adjusted policies, particularly those with exponential growth, not only converge more stably but also do so at a reduced computational cost. This work has broad implications for the development of AI models, especially in domains where the generation of synthetic data is critical for enhancing model capabilities without the extensive need for expensive or hard-to-obtain real-world data. The practical applications of this research extend to fields such as autonomous driving, medical image analysis, and any other area where AI must perform with high stability and efficiency under constraints of limited labeled data. This study thus provides a foundation for more sustainable AI development practices that prioritize both performance and computational efficiency.
@@ Appendix A Theoretic Intuition
We show the proof of Theorem 3.1.
Proof of Theorem 3.1.
@@
$p(\theta^{(T)}\mid\theta^{(0)},\pi)=\mathcal{N}\left(\frac{\theta^{(0)}}{(1+\sigma^{2}/\kappa^{2})^{T}},\sigma^{2}\sum_{t=0}^{T-1}\frac{1}{n_{t}(1+\sigma^{2}/\kappa^{2})^{2(T-t)-1}}\right)$
Proof of Lemma A.1.
Suppose that there are $N_{t}$ data $\{x_{i}\}_{i=1}^{N_{t}}$ generated from the $t$-th step generator which satisfies that the selected sample size is $n_{t}$. Let $b_{i}\sim\operatorname{Bernoulli}(\mathcal{R}(x_{i}))$ as a selecting sign where $b_{i}=1$ indicates that $x_{i}$ is selected and placed into $\bar{D}_{t}$, otherwise $b_{i}=0$ indicates that $x_{i}\notin\bar{D}_{t}$.
We first consider one-step updating. We have the distribution of a data $x_{i}$ conditioned on $x_{i}\in\bar{D}_{t}$,
$p\left(x_{i}\mid x_{i}\in\bar{D}_{t}\right)$ $=p\left(x_{i}\mid b_{i}=1\right)$
$\propto\mathcal{R}(x_{i})\cdot\mathcal{N}\left(x_{i}\mid\theta^{(t)},\sigma^{2}\right)$
@@
$\theta^{(t+1)}={\arg\max}_{\theta}\prod_{x\in\bar{D}_{t}}\mathcal{N}(x;\theta,\sigma^{2})=\frac{1}{n}\sum_{x\in\bar{D}_{t}}x_{i},$
@@
According to Lemma A.1, we have that
$\mathbb{E}_{\theta^{(T)}}\mathbb{E}_{x}\mathcal{R}(x)$ $=\mathbb{E}_{\theta^{(T)}}\mathbb{E}_{x}\exp\left(-\frac{x^{2}}{2\kappa^{2}}\right)$
@@
$=\frac{\kappa}{\sqrt{\sigma^{2}+\kappa^{2}+\sigma_{T}^{2}}}\exp\left(-\frac{\mu_{T}^{2}}{2(\sigma^{2}+\kappa^{2}+\sigma_{T}^{2})}\right)$
@@
Reindexing or shifting $t$ will complete the proof. ∎

### v2

@@ 5 Experiments
We conduct two experiments to evaluate our schemes: image denoising with DPMs, and mathematical reasoning with LLMs, covering both image and text domains. The setups and implementation details are summarized in Table˜1.
In these settings, we emphasize that our goal is not to propose a novel task-solving method or generation technique. Rather, we focus on optimizing the allocation of generation and training costs in iterative learning with synthetic data—an objective that is method-agnostic and can be applied to any generation approach. Here, we validate our theoretical results from Section˜4 and conduct a performance comparison among exponential, linear, and constant schemes on image data (using DPMs) and text data (with LLMs).
5.1 Image Denoising with Diffusion Models
We consider applying iterative learning to the image denoising task, a representative problem in image processing, where the goal is to recover $\bm{x}$ from a noisy observation $\bm{y}=\bm{x}+\bm{\eta}$ with an unknown Gaussian noise $\bm{\eta}$.
@@ 5.1 Image Denoising with Diffusion Models
We choose diffusion probabilistic models (DPMs) (Ho et al., 2020; Song et al., 2021) as denoisers due to their intrinsic design, which aligns closely with the denoising process (Xie et al., 2023) since the reverse process of a DPM gradually estimates and removes this noise using a trained network, effectively performing denoising. Specifically, we fine-tune a pre-trained diffusion model [^1] https://huggingface.co/google/ddpm-cifar10-32 for denoising on the MNIST (Deng, 2012) dataset, and readers may refer to Appendix˜C for more details.
We follow the iterative learning framework to annotate synthetic “clean” images to improve the denoising process. In each iteration, the diffusion model first generates synthetic denoised images $\hat{\bm{x}}$ from the noisy observations. These are then filtered with a reward model linear to the Peak Signal-to-Noise Ratio (PSNR) values compared with the real images. The filtered synthetic data are then used to fine-tune the model through a standard noise-prediction training procedure, iteratively improving its denoising capability. Although an oracle reward model may not be practical in many real-world scenarios, its use here helps to validate the generality of our theoretical results. Furthermore, existing models or correction functions with prior knowledge could be utilized as the reward, as demonstrated in Gillman et al. (2024).
Figure 3: Empirical results of image denoising. The figure shows the average PSNR of the generated denoised images as a function of computational cost. Computational cost is measured in floating-point operations (FLOPs) during iterative learning, including both generation and training. The training batch size is set to $B=640$. We adopt $n_{t}=1.1^{t}\cdot B$ and $n_{t}=1.05^{t}\cdot B$ as the exponential growth policies for $s=10$ and $s=20$, respectively.
Implementations.
We perform denoising for two noise levels corresponding to the diffusion step $s=10$ and $20$, where the PSNR of noisy images are 33.25 and 28.40, respectively. A summary of the algorithm is provided in Algorithm˜2. We consider three constant and three linear policies, each with small, medium, and large $n_{0}$. For the exponential policy, we have performed a small hyperparameter search over $n_{0}$ and $u$, selecting good configurations for two settings. The complete results of the hyperparameter search are presented as an ablation study on different exponential schemes in Section˜F.2, demonstrating its robustness. The configurations for all schemes are presented in the caption and legend of Figure˜3.
@@ Results.
We summarize the final performance of different schemes and baselines in Table˜2 and show the trade-off between performance and computational cost in Figure˜3. We observe that the exponential policy yields the best performance for $s=10$, and performs comparable to the best linear policy for $s=20$. This observation demonstrates both the benefits of an increasing policy (Theorem˜4.2) and the robustness of the exponential growth policy (Theorem˜4.4). In contrast, the constant policy only matches the exponential policy in the early stages of training, but soon reaches a performance plateau and eventually decays due to overfitting, corroborating Theorem˜4.1.
Table 2: Comparison of each policy on the image denoising and math reasoning tasks. The final accuracy are determined using a held-out validation set. We use boldface for the best result while the underline for the second best. For the exponential policies, we only exhibit a single set of configurations for the comparison. We plot for all the policies in Figure˜3 and Figure˜4, respectively.
(a) Image denoising
Policy $n_{t}$ $s=10$ $s=20$
@@
exponential 44.84 40.14
(b) Math reasoning
$n_{t}$ symbolic p1 p2
N/A 55.60 34.98 15.72
$10\cdot B$ 60.28 38.50 16.93
@@
65.66 47.26 20.65
5.2 Math Reasoning with LLMs
Now we move to natural language processing. Iterative learning with synthetic data is also an emerging method for LLMs. Previous research (Zelikman et al., 2022) leverages the model to generate synthetic data, which is then filtered using a reward model or accuracy metrics and subsequently utilized to fine-tune the model. This framework is employed in various settings, including instruction tuning (Touvron et al., 2023), alignment (Dong et al., 2023), and recent efforts to enhance reasoning capabilities through chain-of-thought (CoT) approaches (Hosseini et al., 2024; Lehnert et al., 2024; Singh et al., 2025).
Setup.
We select mathematical reasoning to evaluate the effectiveness of our strategy in iterative learning. Let $q$ and $a$ represent the question and its corresponding answer, respectively. We use the LLM to generate synthetic solutions with CoT and in-context examples. These synthetic data are filtered according to the correctness of the final generated answer $\hat{a}$, defined as the reward $\mathcal{R}(\hat{a})=\mathbf{1}(\hat{a}=a)$. The filtered data are used to further train the model in an iterative process, without filtering based on the correctness of CoT itself. In summary, synthetic data with the correct final solutions are generated and used for training in each iteration, closely mimicking the STaR method (Zelikman et al., 2022).
Implementations.
Many contemporary LLMs are known to overfit to GSM8k (Cobbe et al., 2021) or exhibit a certain degree of data leakage (Zhang et al., 2024; Mirzadeh et al., 2025), where even minor changes to numbers or names can result in significant performance degradation. This issue renders the evaluation results on GSM8k unreliable. To address this, we regenerate three math datasets—GSM_symbolic, GSM_p1, and GSM_p2—using symbolic math templates from Mirzadeh et al. (2025), one of the most outstanding papers on evaluating math reasoning. Each dataset consists of questions paired with their correct answers. GSM_symbolic retains a similar difficulty level to GSM8k, while GSM_p1 and GSM_p2 introduce one and two additional clauses per question, respectively, making them progressively more challenging. Together, these datasets complete a robust evaluation set.
A summary of the algorithm is provided in Algorithm˜3. Similarly to image denoising, we consider three constant and three linear policies, each with small, medium, and large $n_{0}$. For the exponential scheme, we have selected one configuration and left the full ablation results in Section˜F.3. The configurations for all schemes are detailed in Figure˜4 and its caption. We also include the performance of the pre-trained model as a baseline.
Figure 4: Empirical results of math reasoning. The figure shows the accuracies with respect to the computational cost, measured in FLOPs as well. The training batch size is set to $B=256$. We adopt $n_{t}=10\cdot 2^{t}\cdot B$ as the exponential policy. We set the temperature to 0.3 during the generation step and include the results for a temperature of 0.7 in Section˜F.4, which lead to the same conclusions.
@@
Figure˜4 shows the performance versus computational cost across the three datasets under iterative training, while Table˜2 summarizes the final performance for all policies. Across all three difficulty levels, the exponential growth policy demonstrates the steadiest improvement compared to other schemes. Although the linear policy ranks second overall, it fails to match the exponential growth policy when the training budget is large. In particular, with the exponential growth policy, iterative learning increases accuracy on GSM_p1 from 35% to 47% and on GSM_p2 from 16% to 21%. Based on these results, we recommend the exponential scheme for iterative learning.
@@ 6 Conclusion
In this work, we take the first step toward understanding how to design optimal budget allocation policies in iterative bootstrapping for supervised fine-tuning. Through a combination of theoretical analysis and empirical validations, we demonstrate that constant policies are insufficient for achieving convergence, whereas increasing policies offer a more effective alternative. Among increasing policies, the exponential growth policy emerges as a robust and efficient method.
@@ Appendix A Theoretic Intuition
A.1 Discussion on the Gaussian Setting
In the Gaussian case, we use a simplified toy example that omits the generation cost only to provide a clear, introductory setup for the reader. In our complete framework (see Section˜4), however, we assume a generation cost for standard estimators and incorporate this cost into the objective using the hyperparameter $c_{g}$. This hyperparameter offers flexibility: setting $c_{g}=0$ corresponds to free generation, while a large $c_{g}$ reflects expensive generation. We also provide general theorems that hold for all values of $c_{g}$, and explain how the Gaussian case satisfies all assumptions of the general theorems in Section˜B.2. Thus, the Gaussian case can be viewed as a special instance, and the theoretical results still remain valid regardless of whether the sampling is free or costly.
A.2 Proof of Theorem˜3.1
@@ A.1 Discussion on the Gaussian Setting
$p(\theta^{(T)}\mid\theta^{(0)},\pi)=\mathcal{N}\left(\frac{\theta^{(0)}}{(1+\sigma^{2}/\kappa^{2})^{T}},\sigma^{2}\sum_{t=0}^{T-1}\frac{1}{n_{t}(1+\sigma^{2}/\kappa^{2})^{2(T-t)-1}}\right).$
Proof.
Suppose that there are $N_{t}$ data $\{x_{i}\}_{i=1}^{N_{t}}$ generated from the $t$-th step generator such that the selected sample size is $n_{t}$. Let $b_{i}\sim\operatorname{Bernoulli}(\mathcal{R}(x_{i}))$ as a selecting sign where $b_{i}=1$ indicates that $x_{i}$ is selected and placed into $D_{t}$, otherwise $b_{i}=0$ indicates that $x_{i}\notin D_{t}$.
We first consider one-step updating. We have the distribution of a data $x_{i}$ conditioned on $x_{i}\in D_{t}$,
$p\left(x_{i}\mid x_{i}\in D_{t}\right)$ $=p\left(x_{i}\mid b_{i}=1\right)$
$\propto\mathcal{R}(x_{i})\cdot\mathcal{N}\left(x_{i};\theta^{(t)},\sigma^{2}\right)$
@@
$\theta^{(t+1)}={\arg\max}_{\theta}\prod_{x\in D_{t}}\mathcal{N}(x;\theta,\sigma^{2})=\frac{1}{n}\sum_{x\in D_{t}}x_{i},$
@@
According to Lemma˜A.1, we have that
$\mathbb{E}_{\theta^{(T)}}\mathbb{E}_{x}[\mathcal{R}(x)]$ $=\mathbb{E}_{\theta^{(T)}}\mathbb{E}_{x}\left[\exp\left(-\frac{x^{2}}{2\kappa^{2}}\right)\right]$
@@
$=\frac{\kappa}{\sqrt{\sigma^{2}+\kappa^{2}+\sigma_{T}^{2}}}\exp\left(-\frac{\mu_{T}^{2}}{2(\sigma^{2}+\kappa^{2}+\sigma_{T}^{2})}\right).$
@@
Reindexing or shifting $t$ will complete the proof.